A decrease in the price of a product from R50,00 to R40,00 causes the quantity demanded to
increase from 2 500 to 3 000 units. Using the arc elasticity of demand, the price elasticity of
1
Expert's answer
2018-04-09T02:51:15-0400
E_d^p=((Q2-Q1)/((Q2+Q1)/2))/((P2-P1)/((P2+P1)/2))=((3000-2500)/((3000+2500)/2))/((40-50)/((40+50)/2))=0.18/(-0.22)=-0.82